### v1

# Score Change of Variables

Abstract
We derive a general change of variables formula for score functions, showing that for a smooth, invertible transformation $\mathbf{y}=\phi(\mathbf{x})$, the transformed score function $\nabla_{\mathbf{y}}\log q(\mathbf{y})$ can be expressed directly in terms of $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Using this result, we develop two applications: First, we establish a reverse-time Itô lemma for score-based diffusion models, allowing the use of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ to reverse an SDE in the transformed space without directly learning $\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})$. This approach enables training diffusion models in one space but sampling in another, effectively decoupling the forward and reverse processes. Second, we introduce generalized sliced score matching, extending traditional sliced score matching from linear projections to arbitrary smooth transformations. This provides greater flexibility in high-dimensional density estimation. We demonstrate these theoretical advances through applications to diffusion on the probability simplex and empirically compare our generalized score matching approach against traditional sliced score matching methods.
1 Introduction
The change of variables formula is a cornerstone of mathematical analysis, crucial for understanding how quantities transform under coordinate mappings [10, 13]. Its applications span numerous fields, from physics to probability theory. In probability theory, this formula describes how probability density functions (PDFs) relate under a differentiable invertible transformation $\phi$: if a random variable $\mathbf{X}$ follows a density $p(\mathbf{x})$, then $\mathbf{Y}=\phi(\mathbf{X})$ follows a density
$q(\mathbf{y})=p(\phi^{-1}(\mathbf{y}))\left|\det\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right|,$ (1)
where $\mathbf{J}_{\phi^{-1}}$ denotes the Jacobian matrix of the inverse transformation $\phi^{-1}$.
Recently, the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ has gained significant attention in machine learning, playing a pivotal role in diffusion models for generative modeling, statistical inference, and density estimation [25, 14, 26, 9]. State-of-the-art diffusion models, such as those in [7], have demonstrated remarkable success in generating high-fidelity samples by leveraging the score function. Similarly, score matching techniques [16, 26, 23] offer an efficient way to estimate the score directly from data, bypassing the need for explicit density estimation and enabling scalability to high-dimensional problems.
Despite the increasing prominence of the score function, its transformation properties under coordinate changes remain widely unexplored. In this work, we address this limitation by deriving a change of variables formula for score functions. Our formula expresses the score function in the transformed space, $\nabla_{\mathbf{y}}\log q(\mathbf{y})$, directly in terms of the score function in the original space, $\nabla_{\mathbf{x}}\log p(\mathbf{x})$, and derivatives of the transformation $\phi$. This result provides a mathematical tool for adapting score-based methods to transformed spaces, analogous to how the classical change of variables formula facilitates the transformation of probability densities.
Furthermore, we explore the implications of our score function transformation formula in the context of diffusion models and score matching. We derive a reverse-time Itô lemma that enables efficient sampling in transformed spaces using score functions learned in the original space, effectively decoupling the forward and reverse processes. This decoupling allows for greater flexibility in designing diffusion models tailored to specific geometries or constraints. Additionally, we introduce Generalized Sliced Score Matching (GSSM), which extends the idea of linear projections in sliced score matching. Similar to how sliced score matching reduces dimensionality using a linear projection, GSSM achieves this by projecting onto a one-dimensional space defined by the gradient of a general smooth function, providing a non-linear generalization of the original concept. This generalization expands the applicability of sliced score matching, allowing for more flexible and potentially more accurate score estimation in diverse settings.
Our main contributions are:
  1. 1.
    Score Function Change of Variables: We establish a general and explicit formula for transforming score functions under smooth invertible mappings.
  2. 2.
    Reverse-Time Itô Lemma: We develop a reverse-time Itô lemma that allows sampling in transformed spaces by leveraging score functions learned in the original space, enabling more flexible diffusion models.
  3. 3.
    Generalized Sliced Score Matching: We extend sliced score matching [23] by generalizing the linear projection in the original method to the gradient of an arbitrary smooth function from $\mathbb{R}^{n}\text{ to }\mathbb{R}$.
By providing these theoretical contributions, our work paves the way for a deeper understanding and broader application of score-based methods in various machine learning domains. The ability to seamlessly transform score functions across coordinate systems opens up new possibilities for developing more sophisticated and effective generative models and inference algorithms.
2 Main Result
We present a change of variables formula for score functions under smooth, bijective mappings, beginning with the necessary assumptions.
2.1 Assumptions
Assumption 1 (Bijectivity and Smoothness in $\mathbb{R}^{n}$).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, and let $\mathbf{X}:\Omega\to\mathbb{R}^{n}$ be a random vector with probability density function $p$. Assume that $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ is well-defined on $\mathbb{R}^{n}$ and that $\phi:U\to U_{1}$, with $U,U_{1}\subset\mathbb{R}^{n}$, is a bijective and twice continuously differentiable mapping. Let $q(\mathbf{y})$ denote the probability density function of the transformed variable $\mathbf{Y}=\phi(\mathbf{X})$.
2.2 Score Change of Variables
Theorem 1 (Score Change of Variables: $\mathbb{R}^{n}\to\mathbb{R}^{n}$).
Suppose Assumption 1 holds. Then, the score function $\nabla_{\mathbf{y}}\log q(\mathbf{y})$ can be expressed as:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x}))^{\top}\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}.$
Remark 1.
Throughout this paper:
  • $\mathbf{J}_{\phi}$ denotes the Jacobian matrix of the transformation $\phi$.
  • For any matrix-valued function $\mathbf{A}(\mathbf{x}):\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$, we define its divergence row-wise:
    $\nabla_{\mathbf{x}}\cdot\mathbf{A}(\mathbf{x}):=\left[\nabla_{\mathbf{x}}\cdot\mathbf{A}_{i}(\mathbf{x})\right]_{i=1}^{n}\in\mathbb{R}^{n},$
    where $\mathbf{A}_{i}$ is the $i$-th row of $\mathbf{A}$.
Proof.
Starting from the change of variables formula for densities, $q(\mathbf{y})=p(\phi^{-1}(\mathbf{y}))|\det(\mathbf{J}_{\phi^{-1}}(\mathbf{y}))|$, we take the gradient of the log-density and apply the chain rule and the logarithmic derivative of the determinant. Detailed calculations are in appendix A. ∎
2.3 Special Cases
The multidimensional result also covers important special cases, presented as corollaries.
Corollary 1 (Score Change of Variables: $\mathbb{R}\to\mathbb{R}$).
Let $X:\Omega\to\mathbb{R}$ be a random variable with probability density function $p$. Assume that $\nabla_{x}\log p(x)$ is well-defined on $\mathbb{R}$ and that $\phi:U\to U_{1}$, with $U,U_{1}\subset\mathbb{R}$, is a bijective and twice continuously differentiable mapping. Let $q(y)$ denote the probability density function of the transformed variable $Y=\phi(X)$. Then:
$\nabla_{y}\log q(y)=(\phi^{-1})^{\prime}(y)\nabla_{x}\log p(\phi^{-1}(y))+\frac{(\phi^{-1})^{\prime\prime}(y)}{(\phi^{-1})^{\prime}(y)},$
or, equivalently,
$\nabla_{y}\log q(y)=\frac{1}{(\phi^{\prime}(x))^{2}}\left(\phi^{\prime}(x)\nabla_{x}\log p(x)-\phi^{\prime\prime}(x)\right)\bigg{|}_{x=\phi^{-1}(y)}.$
Corollary 2 (Score Change of Variables: $\mathbb{R}^{n}\to\mathbb{R}$).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, and let $\mathbf{X}:\Omega\to\mathbb{R}^{n}$ have probability density function $p$. Let $v:U\to U_{1}$, where $U\subset\mathbb{R}^{n}$ and $U_{1}\subset\mathbb{R}$, be a twice continuously differentiable mapping. Assume that $\|\nabla v(\mathbf{x})\|_{2}\neq 0$ for all $\mathbf{x}\in U$. Let $q(y)$ be the density of $Y=v(\mathbf{X})\in\mathbb{R}$. Then:
$\nabla_{y}\log q(y)=\frac{1}{\|\nabla v(\mathbf{x})\|_{2}^{2}}\left(\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})-\Delta v(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)\right).$
Proof.
The proof constructs a sequence of transformations involving $v$, applies Theorem 1 to each step, and then combines the results. See Appendix A for the full derivation. ∎
3 Applications
We present two applications of the score change of variables: a reverse-time Itô lemma for transformed diffusion processes and a generalized sliced score matching method (Section 3.2). Both applications leverage the score change of variables formula for more flexible training and sampling.
3.1 Reverse-Time Itô Lemma
Score-based diffusion models [7, 15, 21] generate samples by reversing a forward SDE that progressively adds noise to the data. This reversal relies on the score function of the marginal distributions, $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. Given a forward SDE:
$\mathrm{d}\mathbf{X}=\mathbf{f}(\mathbf{X},t)\,\mathrm{d}t+\mathbf{G}(\mathbf{X},t)\,\mathrm{d}\mathbf{W}(t),$ (2)
where $\mathbf{W}(t)$ is a standard Wiener process, the corresponding reverse-time SDE [1, 25] is:
$\mathrm{d}\mathbf{X}$ $=\left[\mathbf{f}(\mathbf{X},t)-\mathbf{G}(\mathbf{X},t)\mathbf{G}(\mathbf{X},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{X})-\nabla_{\mathbf{x}}\cdot\left(\mathbf{G}(\mathbf{X},t)\mathbf{G}(\mathbf{X},t)^{\top}\right)\right]\mathrm{d}t$ (3)
$\quad+\mathbf{G}(\mathbf{X},t)\,\mathrm{d}\bar{\mathbf{W}}(t),$ (4)
where $\bar{\mathbf{W}}(t)$ is a Wiener process in reverse time. This reverse process allows us to recover the original data distribution by starting from noise.
In many scenarios, we prefer to perform this generative diffusion process in a transformed space, for instance when data lies on a manifold [18, 6]. Consider a bijective, twice continuously differentiable transformation $\phi:\mathbb{R}^{n}\times[0,T]\to\mathbb{R}^{n}$. The transformed process $\mathbf{Y}(t)=\phi(\mathbf{X}(t),t)$ for $t\in[0,T]$ satisfies a new SDE [17]:
$\mathrm{d}\mathbf{Y}=\bar{\mathbf{f}}(\mathbf{Y},t)\,\mathrm{d}t+\bar{\mathbf{G}}(\mathbf{Y},t)\,\mathrm{d}\mathbf{W}(t),$ (5)
where the transformed drift $\bar{\mathbf{f}}$ and diffusion $\bar{\mathbf{G}}$ are obtained via Itô’s Lemma:
$\bar{\mathbf{f}}(\mathbf{y},t)$ $=\left(\frac{\partial\phi(\mathbf{x},t)}{\partial t}+\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{f}(\mathbf{x},t)+\frac{1}{2}\text{Tr}\left[\mathbf{G}(\mathbf{x},t)^{\top}\mathbf{H}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\right]\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)},$
$\bar{\mathbf{G}}(\mathbf{y},t)$ $=\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)},$
and $\mathbf{H}_{\phi}$ is the Hessian tensor of $\phi$ with respect to $\mathbf{x}$.
Applying our score change of variables formula leads to the main result for transformed diffusion processes:
Corollary 3 (Reverse-Time Itô Lemma).
Let $\phi:\mathbb{R}^{n}\times[0,\infty)\to\mathbb{R}^{n}$ be a bijective and twice-differentiable transformation for each $t$. Let $p_{t}(\mathbf{x})$ and $q_{t}(\mathbf{y})$ be the probability density functions at time $t$ of $\mathbf{X}$ and $\mathbf{Y}=\phi(\mathbf{X},t)$, respectively. The reverse-time SDE of $\mathbf{Y}$ is:
$\mathrm{d}\mathbf{Y}$ $=\left[\bar{\mathbf{f}}(\mathbf{Y},t)-\nabla_{\mathbf{y}}\cdot\bigl{(}\bar{\mathbf{G}}(\mathbf{Y},t)\bar{\mathbf{G}}(\mathbf{Y},t)^{\top}\bigr{)}\right.$
$\quad\left.-\bar{\mathbf{G}}(\mathbf{Y},t)\bar{\mathbf{G}}(\mathbf{Y},t)^{\top}\left(\mathbf{J}_{\phi^{-1}}(\mathbf{Y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\phi^{-1}(\mathbf{Y},t))\right.\right.$
$\quad\quad\left.\left.+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t),t)^{\top}\right)\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)}\right)\right]\mathrm{d}t+\bar{\mathbf{G}}(\mathbf{Y},t)\mathrm{d}\bar{\mathbf{W}}(t).$ (6)
This result enables sampling in the transformed space while using score functions learned in the original space. In practice, we approximate $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ with a neural network $s_{\theta}(\mathbf{x},t)$ trained via denoising score matching [26]. To account for the transformation, we modify the training loss to incorporate appropriate Jacobian weights (see Appendix A.4 for details):
$L(\theta)=\mathbb{E}_{t}\left[\lambda(t)\,\mathbb{E}_{\mathbf{X}(0)}\,\mathbb{E}_{\mathbf{X}|\mathbf{X}(0)}\left\|\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t),t)^{\top}\left(s_{\theta}(\mathbf{X},t)-\nabla_{\mathbf{x}}\log p_{0t}(\mathbf{X}|\mathbf{X}(0))\right)\right\|^{2}\right]$ (7)
This weighted objective ensures that errors in the score approximation transform appropriately under $\phi$.
3.2 Transformed Score Matching
While the previous section applied Theorem 1 to continuous-time processes, we now generalize sliced score matching using Corollary 2. This Generalized Sliced Score Matching (GSSM) framework offers greater flexibility in data projection while retaining computational efficiency. We first briefly review score matching and sliced score matching, and then introduce GSSM.
3.2.1 Background: Score Matching
Score matching [16] estimates the score function $\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})$ of a probability density $p_{d}(\mathbf{x})$ from samples $\mathbf{x}\sim p_{d}$, without requiring an explicit density form. Consider the loss for a score model $s_{\theta}(\mathbf{x})$:
$\mathcal{L}_{\text{SM}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}}\left[\|s_{\theta}(\mathbf{X})-\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\|^{2}\right].$ (8)
Since $\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})$ is unknown, we cannot compute this directly. Using integration by parts and suitable boundary conditions, the loss can be rewritten without directly involving $\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})$:
$\mathcal{L}_{\text{SM}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}}\left[\|s_{\theta}(\mathbf{X})\|^{2}\right]+\mathbb{E}_{p_{d}}\left[\nabla_{\mathbf{x}}\cdot s_{\theta}(\mathbf{X})\right].$ (9)
3.2.2 Sliced Score Matching
In high-dimensional settings, evaluating $\mathbb{E}_{p_{d}}[\nabla_{\mathbf{x}}\cdot s_{\theta}(\mathbf{X})]$ can be prohibitively expensive. Sliced score matching [23] addresses this by projecting onto random one-dimensional subspaces. Specifically, it uses random vectors $\mathbf{v}\in\mathbb{R}^{n}$ drawn from a distribution $p_{\mathbf{v}}$, where $\mathbb{E}_{p_{\mathbf{v}}}[\|\mathbf{v}\|^{2}]<\infty$ and $\mathbb{E}_{p_{\mathbf{v}}}[\mathbf{v}\mathbf{v}^{\top}]\succ 0$. Under these conditions, the loss becomes:
$\mathcal{L}_{\text{SSM}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{d}}\mathbb{E}_{p_{\mathbf{v}}}\left[\left(\mathbf{v}^{\top}s_{\theta}(\mathbf{X})\right)^{2}\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{\mathbf{v}}}\left[\mathbf{v}^{\top}\nabla_{\mathbf{x}}\left(\mathbf{v}^{\top}s_{\theta}(\mathbf{X})\right)\right].$ (10)
This objective is equivalent to the original score matching loss up to a constant, yet is more tractable computationally in high dimensions.
3.2.3 Generalized Sliced Score Matching
We now introduce Generalized Sliced Score Matching (GSSM), which extends sliced score matching to use arbitrary smooth transformations. Let $v:\mathbb{R}^{n}\to\mathbb{R}$ be a twice continuously differentiable random function drawn from an independent distribution $p_{v}$. Using Corollary 2, we derive:
$\mathcal{L}_{\text{GSSM}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\left(\nabla_{\mathbf{x}}v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)^{2}\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[(\nabla_{\mathbf{x}}v(\mathbf{X}))^{\top}\nabla_{\mathbf{x}}\bigl{(}s_{\theta}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\bigr{)}\right]$
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\nabla_{\mathbf{x}}v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right].$ (11)
The extra terms compared to SSM account for the nonlinearity of $v$ via its Hessian $\mathbf{H}_{v}$ and Laplacian $\Delta v$.
When $v(\mathbf{x})=\mathbf{v}^{\top}\mathbf{x}$ is linear, we have $\nabla_{\mathbf{x}}v(\mathbf{x})=\mathbf{v}$, $\mathbf{H}_{v}=0$, and $\Delta v=0$, reducing GSSM to the original sliced score matching loss (10). Thus, GSSM strictly generalizes sliced score matching, providing more flexibility and potentially improved performance in complex, high-dimensional scenarios.
3.2.4 Derivation Outline
To derive GSSM, we first apply standard score matching in the transformed space $y=v(\mathbf{x})$. Then, using Corollary 2, we substitute the transformed score back into the original space. Integration by parts under appropriate conditions (see Appendix B) yields Equation 3.2.3. The complete derivation is in Appendix B.
4 Examples
4.1 Diffusion on the Probability Simplex for Chess Positions
We demonstrate our reverse-time Itô lemma by generating chess positions through diffusion on a probability simplex. Prior work has explored diffusion on the simplex by training directly in a constrained space [12], but our approach enables training in an unconstrained space with a Gaussian score model while still sampling in the simplex.
4.1.1 Representing Chess Positions
A chess position can be represented as a point in a 13-dimensional probability simplex. We focus on the projected simplex defined as:
$\{\mathbf{y}\in[0,1]^{12}\mid\sum_{i=1}^{12}y_{i}\leq 1\}.$
Here, each element $\mathbf{y}$ specifies probabilities for 12 possible piece types (6 for White and 6 for Black) that could occupy a single square on the chessboard. The probability of the square being empty is implicitly given by $1-\sum_{i=1}^{12}y_{i}$. Extending this construction to all 64 squares, we represent a full chess position as a collection of such probability vectors across the board.
4.1.2 Transformation Between Spaces
We employ the additive logistic transformation [3] and its inverse to map between $\mathbb{R}^{12}$ and the projected simplex. For $i=1,\dots,12$:
$\phi_{i}(\mathbf{x})$ $=\frac{e^{x_{i}}}{1+\sum_{j=1}^{12}e^{x_{j}}},$ (12)
$\psi_{i}(\mathbf{p})$ $=\log\left(\frac{p_{i}}{1-\sum_{j=1}^{12}p_{j}}\right).$ (13)
This mapping allows us to freely train a score-based model in an unconstrained Euclidean space, then transform and sample in the simplex domain where the probabilities must sum to at most one.
4.1.3 Training and Sampling
We train a score-based model $s_{\theta}(\mathbf{x},t)$ in $\mathbb{R}^{12}$. Chess positions are represented with a softened one-hot encoding to avoid infinite values when applying the inverse transformation $\psi(\cdot)$. The score model is trained using a Variance Preserving (VP) SDE [25]:
$\mathrm{d}\mathbf{X}=-\frac{1}{2}\beta(t)\mathbf{X}\,\mathrm{d}t+\sqrt{\beta(t)}\,\mathrm{d}\mathbf{W},$ (14)
where $\beta(t)$ is the noise schedule. During inference, we apply our reverse-time Itô lemma to perform diffusion directly in the projected simplex. By leveraging the learned scores in the unconstrained space and the known transformation $\phi$, we obtain the appropriate transformed SDE coefficients (derived in Appendix E) and sample valid chess positions on the simplex.
4.1.4 Controlling Piece Density
The transformed drift term, $\bar{\mathbf{f}}(\mathbf{Y},t)$, interacts with the simplex geometry to influence piece distributions. To gain intuitive control, we introduce a scaling factor $w$:
$\tilde{\mathbf{f}}(\mathbf{Y},t)=w\bar{\mathbf{f}}(\mathbf{Y},t).$ (15)
Increasing $w$ biases the distribution toward fewer pieces (increasing empty squares), while decreasing it encourages more pieces. This simple scalar parameter provides interpretable geometric control over the final configurations, allowing practitioners to shape the sampling process easily. The generated positions in Figure 5 demonstrate how varying $w$ changes piece density without losing consistency in the underlying chess position representation.
Figure 1: Generated positions with $w=0.9$.
Figure 2: Generated positions with $w=1.0$.
Figure 3: Generated positions with $w=1.1$.
Figure 4: Generated positions with $w=1.2$.
Figure 5: Generated chess positions with different values of $w$. Increasing $w$ reduces the number of pieces on the board, illustrating the intuitive geometric control enabled by sampling in the transformed simplex space.
4.2 Density Estimation with Generalized Sliced Score Matching
We evaluate Generalized Sliced Score Matching (GSSM) on density estimation tasks using deep kernel exponential families (DKEFs), comparing its performance against traditional score matching (SM) and sliced score matching (SSM) [23]. We also introduce and evaluate a variance-reduced version of GSSM (GSSM-VR).
4.2.1 Choice of Transformation
For GSSM, we use a quadratic form as our nonlinear transformation:
$v(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}\mathbf{A}\mathbf{x}+\mathbf{b}^{\top}\mathbf{x},$ (16)
where $\mathbf{A}$ is a random symmetric matrix and $\mathbf{b}$ is a random vector. The entries of $\mathbf{A}$ and $\mathbf{b}$ are drawn from zero-mean distributions with specified variances: $\sigma_{1}^{2}$ for diagonal entries of $\mathbf{A}$, $\sigma_{2}^{2}$ for off-diagonal entries, and $\sigma_{3}^{2}$ for entries of $\mathbf{b}$. The gradient of this transformation, which determines the direction of projection in GSSM, is:
$\nabla_{\mathbf{x}}v(\mathbf{x})=\mathbf{A}\mathbf{x}+\mathbf{b}.$ (17)
4.2.2 Variance Reduction
Depending on the chosen transformation, we may reduce the variance of our estimator by analytically integrating out some randomness. For linear transformations (as in standard sliced score matching), integrating out the random directions yields the known variance-reduced SSM objective [23]:
$\mathcal{L}_{\text{SSM-VR}}(s_{\theta})=\mathbb{E}_{p_{d}}\left[\frac{1}{2}\|s_{\theta}(\mathbf{X})\|^{2}+\mathbb{E}_{p_{\mathbf{v}}}\left[\mathbf{v}^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})\mathbf{v}\right]\right].$ (18)
Similarly, for the quadratic transformation, integrating out some of the randomness from $\mathbf{A}$ and $\mathbf{b}$ leads to a variance-reduced GSSM (GSSM-VR) objective:
$\mathcal{L}_{\text{GSSM-VR}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{d}}\Big{[}\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{X})^{2}x_{i}^{2}+\sigma_{2}^{2}\left(\|\mathbf{X}\|^{2}\|s_{\theta}(\mathbf{X})\|^{2}+(s_{\theta}(\mathbf{X})^{\top}\mathbf{X})^{2}\right)+\sigma_{3}^{2}\|s_{\theta}(\mathbf{X})\|^{2}\Big{]}$ (19)
$\quad+\mathbb{E}_{p_{d}}\left[\left(2\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}\right)s_{\theta}(\mathbf{X})^{\top}\mathbf{X}\right]$
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})(\mathbf{A}\mathbf{X}+\mathbf{b})\right].$
The full derivation for the quadratic case is provided in Appendix C.
4.2.3 Score Function Representation in DKEF Models
We follow [23, 27] and use DKEF models to evaluate different score matching methods. A DKEF approximates an unnormalized log density as:
$\log\tilde{p}_{f}(\mathbf{x})=f(\mathbf{x})+\log q_{0}(\mathbf{x}),\quad f(\mathbf{x})=\sum_{l=1}^{L}\alpha_{l}k(\mathbf{x},\mathbf{z}_{l}),$ (20)
where $k$ is a mixture of Gaussian kernels centered at learned inducing points $\mathbf{z}_{l}$. The kernel features are defined through a neural network, and we jointly optimize the network parameters, kernel weights $\alpha_{l}$, and inducing points $\mathbf{z}_{l}$.
Additional details on the DKEF model architecture are provided in Appendix D.
4.2.4 Experimental Setup
We evaluate SM, SSM, GSSM, and their variance-reduced variants on three UCI datasets (Parkinsons, RedWine, WhiteWine) [2], following the training protocol and hyperparameters from [23]. For the quadratic transformation, we set $\sigma_{3}^{2}=1$ and choose the matrix variances as $\sigma_{2}^{2}=1/\sqrt{n}$ and $\sigma_{1}^{2}=2/\sqrt{n}$, where $n$ is the input dimension. The vector $\mathbf{b}$ is sampled from a Rademacher distribution, and entries of $\mathbf{A}$ are sampled from a Gaussian distribution with the given variances. This setup allows efficient sampling of $\mathbf{A}\mathbf{x}$ without explicitly constructing $\mathbf{A}$ (see Appendix C). For SSM-VR and SSM, we sample the random projection vectors $\mathbf{v}$ from a Rademacher distribution, following [23].
4.2.5 Results
Figure 6: Exact Score Matching loss across three UCI datasets; lower loss is better.
Figure 6 shows the score matching loss for different methods across the three datasets. Our variance-reduced GSSM (GSSM-VR) outperforms SSM-VR on RedWine and Parkinsons, while remaining competitive on WhiteWine. GSSM-VR also demonstrates lower variance in its estimates for RedWine and Parkinsons, indicating more stable optimization.
The non-variance-reduced GSSM performs less favorably than SSM under identical conditions, likely due to higher estimator variance impacting the optimization process. As in [23], training was halted after 200 steps of no validation improvement, which proved limiting for GSSM due to its higher variance. Nevertheless, the key contribution of GSSM is its flexibility: extending score matching beyond linear projections allows tailoring the method to data distributions more effectively. Although the quadratic transformations used here are just one example, the results suggest further exploration into choosing more suitable transformation families for specific datasets or applications.
5 Discussion
This work introduces a change of variables formula for score functions, providing a principled approach to adapt score-based methods across different spaces. By enabling transformations that accommodate specific data geometries and constraints, our result increases flexibility in model design and training. We demonstrate this through a reverse-time Itô lemma for transformed diffusion models and a generalized sliced score matching framework. Together, these contributions lay a mathematical foundation for transforming score-based algorithms, opening up new avenues for research and applications.
5.1 Related Work
Our approach complements existing research on distribution transformations, such as normalizing flows [20, 8], which focus on mapping densities rather than their scores. It also connects to recent developments in score-based diffusion models on constrained domains and manifolds [6, 4, 12, 18, 11]. In particular, [4] uses a change of variables for biological sequence generation, and our framework provides a unifying perspective and rigorous analytical tools to understand such transformations more broadly.
5.2 Limitations
A primary limitation is the computational cost of computing Jacobians and potential numerical instabilities when applying complex transformations. Our theoretical analysis ensures the existence of transformed score functions but does not guarantee numerical stability or efficient convergence for sampling. Moreover, the requirement of smooth, well-conditioned transformations restricts the class of applicable mappings. Finally, current variance reduction techniques for GSSM are limited to specific families of transformations.
5.3 Future Directions
Several promising directions can further advance this line of work.
First, while we focused on analytically tractable mappings, exploring more expressive transformations, such as invertible neural networks, could offer greater flexibility and adaptability to complex data geometries. This data-driven approach could lead to significant performance improvements, but also increase computational costs.
Second, deeper theoretical study is needed to understand the interplay between transformations and diffusion processes. Such insights could guide the selection of optimal transformations, challenging the prevailing assumption that the same space should be used for both training and sampling. This may lead to improved sampling efficiency and better-tailored geometries for specific applications.
Third, extending the score change of variables framework to other score-based methods, including variational inference [19] and energy-based models [24], presents a compelling direction for future research and could unlock new capabilities within these frameworks.
In summary, the theoretical foundations established here open the door to more versatile and effective score-based methodologies, offering opportunities for improved performance, stability, and interpretability across a range of applications.
Appendix A Proofs
A.1 Proof of Theorem 1: Score Change of Variables in $\mathbb{R}^{n}$
Proof.
The probability density function $q(\mathbf{y})$ of $\mathbf{Y}$ is related to $p(\mathbf{x})$ by the change of variables formula:
$q(\mathbf{y})=p(\phi^{-1}(\mathbf{y}))\left|\det\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)\right|.$
Taking the gradient of the logarithm of $q(\mathbf{y})$, we obtain:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\nabla_{\mathbf{y}}\log p(\phi^{-1}(\mathbf{y}))+\nabla_{\mathbf{y}}\log\left|\det\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)\right|.$
Applying the chain rule to the first term yields:
$\nabla_{\mathbf{y}}\log p(\phi^{-1}(\mathbf{y}))=\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{\top}\cdot\nabla_{\mathbf{x}}\log p(\mathbf{x})\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}.$
For the second term, we have:
$\nabla_{\mathbf{y}}\log\left|\det\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)\right|$ $=\frac{\nabla_{\mathbf{y}}\det\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)}{\det\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)}$ (21)
$=\text{Tr}\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{-1}\cdot\nabla_{\mathbf{y}}\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right),$ (22)
where in the second line we use Jacobi’s formula for an invertible matrix.
By matrix calculus, we find:
$\text{Tr}\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{-1}\cdot\nabla_{\mathbf{y}}\mathbf{J}_{\phi^{-1}}(\mathbf{y})\right)$ $=\sum_{i=1}^{n}\mathbf{H}_{\phi_{i}^{-1}}\cdot\frac{\partial\phi}{\partial x_{i}}$ (23)
$=\nabla\cdot\ \mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x}))^{\top},$ (24)
where $\mathbf{H}_{\phi_{i}^{-1}}$ is the Hessian matrix of $\phi_{i}^{-1}$.
Combining both terms, the score function is given by:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{\top}\cdot\nabla_{\mathbf{x}}\log p(\mathbf{x})+\nabla\cdot\ \left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x}))^{\top}\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}.$
Remark 2.
In this proof, we utilize the following result for an invertible function $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ where $f(\mathbf{x})=\mathbf{y}$:
$\nabla_{\mathbf{x}}\log\left|\det\left(\mathbf{J}_{f}(\mathbf{x})\right)\right|=\nabla\cdot\ \mathbf{J}^{\top}_{f}(f^{-1}(\mathbf{y})).$ (25)
We can write:
$\mathbf{J}_{f}(f^{-1}(\mathbf{y}))^{\top}=\left(\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)^{-1}\right)^{\top}=\frac{1}{\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)}\text{Cof}(\mathbf{J}_{f^{-1}}),$ (26)
where $\text{Cof}(\mathbf{J}_{f^{-1}})$ is the cofactor matrix of $\mathbf{J}_{f^{-1}}$.
For all differentiable $f$, we have the fact [10]:
$\nabla_{\mathbf{x}}\cdot\left(\text{Cof}(\mathbf{J}_{f}(\mathbf{x}))\right)=0.$ (27)
Using this, we calculate:
$\nabla_{\mathbf{x}}\log\left|\det\left(\mathbf{J}_{f}(\mathbf{x})\right)\right|$ $=\nabla\cdot\ \mathbf{J}_{f}(f^{-1}(\mathbf{y}))^{\top}$
$=\nabla\cdot\ \frac{1}{\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)}\text{Cof}(\mathbf{J}_{f^{-1}}(\mathbf{y}))$
$=\frac{1}{\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)}\nabla\cdot\ \text{Cof}(\mathbf{J}_{f^{-1}}(\mathbf{y}))+\text{Cof}(\mathbf{J}_{f^{-1}}(\mathbf{y}))\cdot\nabla_{\mathbf{y}}\frac{1}{\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)}$
$=0-\frac{1}{\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)^{2}}\text{Cof}(\mathbf{J}_{f^{-1}}(\mathbf{y}))\cdot\nabla_{\mathbf{y}}\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)$
$=-\mathbf{J}^{\top}_{f}\cdot\nabla_{\mathbf{y}}\log\left|\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)\right|.$
From our above analysis, we have for an invertible and twice-differentiable function $f$:
$\nabla_{\mathbf{x}}\log\left|\det\left(\mathbf{J}_{f}(\mathbf{x})\right)\right|=-\mathbf{J}_{f}^{\top}\cdot\nabla_{\mathbf{y}}\log\left|\det\left(\mathbf{J}_{f^{-1}}(\mathbf{y})\right)\right|.$ (28)
A.2 Proof of Corollary 2: Dimension Reduction $\mathbb{R}^{n}\to\mathbb{R}$
We assume $v(\mathbf{x}):U\to U_{1}$ where $U\subset\mathbb{R}^{n}$ and $U_{1}\subset\mathbb{R}$ is a twice continuously differentiable transformation and that $\|\nabla v(\mathbf{x})\|_{2}$ is non-zero for all $\mathbf{x}$ in $U$. Let $y=v(\mathbf{x})$.
Proof.
Consider the transformation
$\phi(\mathbf{x})=(x_{1},\cdots,x_{i-1},v(\mathbf{x}),x_{i+1},\cdots,x_{n})=\mathbf{z}.$
Suppose $\frac{\partial v(\mathbf{x})}{\partial x_{i}}$ is non-zero. By the change of variables formula, we have
$\nabla_{\mathbf{x}}\log p(\mathbf{x})$ $=\mathbf{J}^{\top}_{\phi(\mathbf{x})}\nabla_{\mathbf{z}}\log q(\mathbf{z})+\nabla_{\mathbf{x}}\log|\det(\mathbf{J}_{\phi}(\mathbf{x}))|\bigg{|}_{\mathbf{z}=\phi(\mathbf{x})}$
$=\mathbf{J}^{\top}_{\phi(\mathbf{x})}\nabla_{\mathbf{z}}\log q(\mathbf{z})+\nabla_{\mathbf{x}}\log\left|\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right|\bigg{|}_{\mathbf{z}=\phi(\mathbf{x})}.$
Then we have
$\frac{\partial}{\partial x_{i}}\log p(\mathbf{x})$ $=\frac{\partial v(\mathbf{x})}{\partial x_{i}}\frac{\partial}{\partial z_{i}}\log q(\mathbf{z})+\frac{\partial}{\partial x_{i}}\log\left|\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right|$
$=\frac{\partial v(\mathbf{x})}{\partial x_{i}}\left(\nabla_{y}\log q(y)+\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)\right)+\frac{\partial}{\partial x_{i}}\log\left|\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right|.$
Rearranging, we see
$\left(\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right)^{2}\nabla_{y}\log q(y)$ $=\frac{\partial v(\mathbf{x})}{\partial x_{i}}\frac{\partial}{\partial x_{i}}\log p(\mathbf{x})$
$-\left(\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)-\frac{\partial^{2}v(\mathbf{x})}{\partial x_{i}^{2}}.$ (29)
Since A.2 is true for all $i$ where $\frac{\partial v(\mathbf{x})}{\partial x_{i}}$ is non-zero and trivially true where $\frac{\partial v(\mathbf{x})}{\partial x_{i}}=0$, we have
$\|\nabla v(\mathbf{x})\|_{2}^{2}\nabla_{y}\log q(y)$ $=\sum_{i=1}^{n}\frac{\partial v(\mathbf{x})}{\partial x_{i}}\frac{\partial}{\partial x_{i}}\log p(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)-\sum_{i=1}^{n}\frac{\partial^{2}v(\mathbf{x})}{\partial x_{i}^{2}}$
$=\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})-\Delta v(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y).$
Finally, assuming $\|\nabla v(\mathbf{x})\|_{2}$ is non-zero we find
$\nabla_{y}\log q(y)$ $=\frac{1}{\|\nabla v(\mathbf{x})\|_{2}^{2}}\left(\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})-\Delta v(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v(\mathbf{x})}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)\right).$
A.3 Additional result: Dimension Expansion $\mathbb{R}\to\mathbb{R}^{n}$
While not used in the main text, we present an additional change of variables formula for completeness.
Theorem 2 (Change of Variables Formula $\mathbb{R}\to\mathbb{R}^{n}$).
Let $v:\mathbb{R}\to\mathbb{R}^{n}$ be a differentiable transformation. Let $p(x)$ be the probability density function of a random variable $X\in\mathbb{R}$, and let $q(\mathbf{y})$ be the probability density function of the transformed random variable $\mathbf{Y}=v(X)\in\mathbb{R}^{n}$. Assume that $\frac{dv_{i}(x)}{dx}$ is non-zero for all $x\in\mathbb{R}$ and for all $i$. Then, the score function $\nabla_{\mathbf{y}}\log q(\mathbf{y})$ of $\mathbf{Y}$ can be expressed as follows:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\frac{1}{\left(\frac{dv(x)}{dx}\right)^{2}}\left(\frac{dv(x)}{dx}\nabla_{x}\log p(x)-\frac{d^{2}v(x)}{dx^{2}}\right),$ (30)
where $\frac{dv(x)}{dx}\in\mathbb{R}^{n}$ is the vector of first derivatives and $\frac{1}{(\frac{dv(x)}{dx})^{2}}$ denotes element-wise division by the squared components of $\frac{dv(x)}{dx}$.
Proof.
Consider the transformation $v:\mathbb{R}\to\mathbb{R}^{n}$ defined by $\mathbf{y}=v(x)$ where $v$ is differentiable, and each component $v_{i}(x)$ of $v$ has a non-zero derivative $\frac{dv_{i}(x)}{dx}$. For each $i$, $v_{i}$ acts essentially as a scalar transformation from $\mathbb{R}\to\mathbb{R}$, making $v_{i}^{-1}$ well defined.
Applying corollary 1 to each component $v_{i}$ independently, we find the score function for each transformed component $y_{i}=v_{i}(x)$ of the random variable $\mathbf{Y}$:
$\nabla_{y_{i}}\log q(y_{i})=\frac{1}{\left(\frac{dv_{i}(x)}{dx}\right)^{2}}\left(\frac{dv_{i}(x)}{dx}\nabla_{x}\log p(x)-\frac{d^{2}v_{i}(x)}{dx^{2}}\right).$
For the multidimensional variable $\mathbf{Y}=v(X)$, the score function $\nabla_{\mathbf{y}}\log q(\mathbf{y})$ can be composed by stacking the gradients of each component $\log q_{i}(y_{i})$, yielding:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\left[\nabla_{y_{1}}\log q(y_{1}),\dots,\nabla_{y_{n}}\log q(y_{n})\right]^{\top}.$
Substituting the expression for each $\nabla_{y_{i}}\log q(y_{i})$ into this vector gives:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\biggl{[}$ $\frac{1}{\left(\frac{dv_{1}(x)}{dx}\right)^{2}}\left(\frac{dv_{1}(x)}{dx}\nabla_{x}\log p(x)-\frac{d^{2}v_{1}(x)}{dx^{2}}\right),$
$\ldots$,… ,
$\frac{1}{\left(\frac{dv_{n}(x)}{dx}\right)^{2}}\left(\frac{dv_{n}(x)}{dx}\nabla_{x}\log p(x)-\frac{d^{2}v_{n}(x)}{dx^{2}}\right)\biggr{]}^{\top}.$
Combining these component-wise expressions, we obtain:
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\frac{1}{\left(\frac{dv(x)}{dx}\right)^{2}}\left(\frac{dv(x)}{dx}\nabla_{x}\log p(x)-\frac{d^{2}v(x)}{dx^{2}}\right),$
where $\frac{dv(x)}{dx}\in\mathbb{R}^{n}$ is the vector of first derivatives $[\frac{dv_{1}(x)}{dx},\ldots,\frac{dv_{n}(x)}{dx}]^{\top}$, $\frac{d^{2}v(x)}{dx^{2}}\in\mathbb{R}^{n}$ is the vector of second derivatives, and all operations (division, multiplication) are performed element-wise. ∎
A.4 Derivation of Weighted Loss Function
We show that minimizing the weighted loss (7) yields a score network that correctly approximates the transformed score. Let $q_{t}(\mathbf{y})$ be the density of the transformed variable $\mathbf{Y}=\phi(\mathbf{X},t)$.
First, recall from Theorem 1 that:
$\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t)t)^{\top}\right)\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)}.$ (31)
Consider the standard denoising score matching objective in the transformed space:
$\mathbb{E}_{t}\left[\lambda(t)\,\mathbb{E}_{\mathbf{Y}(0)}\,\mathbb{E}_{\mathbf{Y}|\mathbf{Y}(0)}\left\|\tilde{s}_{\theta}(\mathbf{Y},t)-\nabla_{\mathbf{y}}\log q_{0t}(\mathbf{Y}|\mathbf{Y}(0))\right\|^{2}\right],$ (32)
where $\tilde{s}_{\theta}$ is a score network in the transformed space and $t$ is sampled uniformly from $[0,1]$.
For this to be equivalent to training in the original space, we want:
$\tilde{s}_{\theta}(\mathbf{y},t)=\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}s_{\theta}(\phi^{-1}(\mathbf{y},t),t)+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right),$ (33)
where $s_{\theta}$ is trained in the original space.
Substituting this relationship and changing variables to $\mathbf{X}$:
$\mathbb{E}_{t}\Big{[}\lambda(t)\,\mathbb{E}_{\mathbf{X}(0)}\,\mathbb{E}_{\mathbf{X}|\mathbf{X}(0)}\Big{\|}\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t))^{\top}s_{\theta}(\mathbf{X},t)+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right)$
$\qquad\qquad-\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t))^{\top}\nabla_{\mathbf{x}}\log p_{0t}(\mathbf{X}|\mathbf{X}(0))-\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right)\Big{\|}^{2}\Big{]}$
$=\mathbb{E}_{t}\Big{[}\lambda(t)\,\mathbb{E}_{\mathbf{X}(0)}\,\mathbb{E}_{\mathbf{X}|\mathbf{X}(0)}\Big{\|}\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t))^{\top}\left(s_{\theta}(\mathbf{X},t)-\nabla_{\mathbf{x}}\log p_{0t}(\mathbf{X}|\mathbf{X}(0))\right)\Big{\|}^{2}\Big{]},$
which is precisely our weighted objective (7).
Therefore, minimizing this weighted loss in the original space is equivalent to training a score network directly in the transformed space. The Jacobian weights ensure that errors in score approximation transform appropriately under $\phi$.
Appendix B Derivation of the Generalized Sliced Score Matching Objective
The derivation of Generalized Sliced Score Matching (GSSM) follows the principles established in the original score matching [16] and sliced score matching [23] papers. We extend these ideas to account for general smooth functions.
Assumption 2 (Integrability Conditions).
For the data distribution $p_{d}(\mathbf{x})$ and score approximation $s_{\theta}(\mathbf{x})$:
  1. 1.
    $\mathbb{E}_{p_{d}}[\|\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\|^{2}]<\infty$
  2. 2.
    $\mathbb{E}_{p_{d}}[\|s_{\theta}(\mathbf{X})\|^{2}]<\infty$
  3. 3.
    For any $v\sim p_{v}$, $\mathbb{E}_{p_{d}}[\|\nabla v(\mathbf{X})\|^{2}]<\infty$ and $\nabla v(\mathbf{X})\neq\mathbf{0}$ almost surely.
Assumption 3 (Boundary Conditions).
For any $v\sim p_{v}$:
$p_{d}(\mathbf{x})(\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x}))\nabla v(\mathbf{x})\to 0\text{ as }\|\mathbf{x}\|\to\infty$ (34)
B.1 Score Matching in Transformed Space
Consider a random vector $\mathbf{X}\in\mathbb{R}^{n}$ with density $p_{d}(\mathbf{x})$. Let $v:\mathbb{R}^{n}\to\mathbb{R}$ be a random twice continuously differentiable function sampled from $p_{v}(v)$, with $\nabla v(\mathbf{x})\neq\mathbf{0}$ almost surely. We denote $q_{d}^{v}$ as the distribution of $Y=v(\mathbf{X})$ when $v$ is fixed. Following the original score matching framework, we minimize:
$\mathcal{L}_{\text{SM}}^{(y)}(\bar{s}_{\theta})=\frac{1}{2}\mathbb{E}_{p_{v}}\left[\mathbb{E}_{q_{d}^{v}}\left[\left(\bar{s}_{\theta}(Y)-\nabla_{y}\log q_{d}^{v}(Y)\right)^{2}\right]\right].$ (35)
From the change of variables formula (Corollary 2):
$\nabla_{y}\log q_{d}^{v}(y)=\frac{1}{\|\nabla v(\mathbf{x})\|^{2}}\left(\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})-R(\mathbf{x})\right)$ (36)
Similarly for the model score:
$\bar{s}_{\theta}(y)=\frac{1}{\|\nabla v(\mathbf{x})\|^{2}}\left(\nabla v(\mathbf{x})^{\top}s_{\theta}(x)-R(\mathbf{x})\right)$ (37)
where $R(\mathbf{x})=\Delta v(\mathbf{x})+\sum_{i=1}^{n}\left(\frac{\partial v}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p_{d}((x_{j})_{j\neq i}\mid y)$.
Substituting these relations into (35):
$\mathcal{L}_{\text{intermediate}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\frac{1}{\|\nabla v(\mathbf{X})\|^{4}}\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})-\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\right)^{2}\right]$ (38)
B.2 Normalization Considerations
The $\|\nabla v(\mathbf{x})\|^{4}$ term in the denominator of (38) disproportionately amplifies the loss when $\|\nabla v(\mathbf{x})\|$ is small. We address this through two alternative approaches.
First, we can weight (38) by $\|\nabla v(\mathbf{x})\|^{2}$, yielding:
$\mathcal{L}_{\text{normalized}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla\tilde{v}(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})-\nabla\tilde{v}(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\right)^{2}\right]$ (39)
where $\nabla\tilde{v}(\mathbf{x})=\frac{\nabla v(\mathbf{x})}{\|\nabla v(\mathbf{x})\|}$. This normalized loss is gradient magnitude-invariant during optimization. However, the normalization introduces computational overhead and makes the variance reduction version of the loss intractable.
As an alternative, we weight (38) by $\|\nabla v(\mathbf{x})\|^{4}$, resulting in:
$\mathcal{L}_{\text{un-normalized}}(s_{\theta})=\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla{v}(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})-\nabla{v}(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\right)^{2}\right]$ (40)
While this loss places greater emphasis on regions where $\|\nabla v(\mathbf{x})\|$ is large, it proves more tractable in practice. We proceed with loss (40) in the following derivations, noting that substituting $\tilde{v}(\mathbf{x})$ for $v(\mathbf{x})$ preserves all subsequent mathematical relationships.
B.3 Eliminating the Dependence on $\nabla_{\mathbf{x}}\log p(\mathbf{x})$
To eliminate $\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})$, we expand the square and apply integration by parts.
First, expand the squared term:
$\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla{v}(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})-\nabla{v}(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})\right)^{2}\right]$
$=\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})\right)^{2}\right]-\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})\right)\left(\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{x})\right)\right]+C$
We simplify the second part of the loss using integration by parts.
Consider the term:
$\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)\left(\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\right)\right]$ (41)
$=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)\left(\nabla v(\mathbf{X})^{\top}\frac{\nabla_{\mathbf{x}}p_{d}(\mathbf{X})}{p_{d}(\mathbf{X})}\right)\right]$ (42)
[…]
We integrate by parts taking in account assumption 3:
[…]
Compute the divergence:
$\nabla_{\mathbf{x}}\cdot\left(\left(\nabla v^{\top}(\mathbf{x})s_{\theta}(\mathbf{x})\right)\nabla v(\mathbf{x})\right)$ $=\left(\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})\right)\Delta v(\mathbf{x})$ (47)
$+\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{H}_{v}(\mathbf{x})\right)\nabla v(\mathbf{x})+\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x})\nabla v(\mathbf{x}).$ (48)
Thus,
$\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)\left(\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\log p_{d}(\mathbf{X})\right)\right]=$
$-\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)\Delta v(\mathbf{X})+s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla v(\mathbf{X})+\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{X}}s_{\theta}\nabla v(\mathbf{X})\right].$
Collecting terms, the loss becomes:
$\mathcal{L}_{\text{GSSM}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)^{2}\right]+\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})\nabla v(\mathbf{X})\right]$
$+\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla v(\mathbf{X})\right]+\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right].$ (49)
Appendix C Quadratic GSSM Details
We choose a quadratic random function:
$v(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}\mathbf{A}\mathbf{x}+\mathbf{b}^{\top}\mathbf{x}$ (50)
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is a random symmetric matrix, and $\mathbf{b}\in\mathbb{R}^{n}$ is a random vector. The random variables are constructed with the following properties:
  • Matrix $\mathbf{A}$:
    • Symmetric: $\mathbf{A}=\mathbf{A}^{\top}$
    • Diagonal entries: $A_{ii}\sim\mathcal{N}(0,\sigma_{1}^{2})$
    • Off-diagonal entries: $A_{ij}=A_{ji}\sim\mathcal{N}(0,\sigma_{2}^{2})$ for $i<j$
    • Independence of entries, except that off-diagonal entries $A_{ij}$ and $A_{ji}$ are dependent due to symmetry
  • Vector $\mathbf{b}$:
    • $\mathbf{b}\sim\mathcal{N}(0,\sigma_{3}^{2}I)$
    • Independent from $\mathbf{A}$
C.1 Quadratic GSSM Loss
The components needed for the GSSM loss are as follows:
$\nabla v(\mathbf{x})$ $=\mathbf{A}\mathbf{x}+\mathbf{b}$ (51)
$\mathbf{H}_{v}(\mathbf{x})$ $=\mathbf{A}$ (52)
$\Delta v(\mathbf{x})$ $=\operatorname{tr}(\mathbf{A})$ (53)
To satisfy the conditions for the GSSM loss 3.2.3, we require $\nabla v(\mathbf{x})\neq\mathbf{0}$. Thus we require $x\neq\mathbf{0}$. For $\mathbf{x}\neq\mathbf{0}$, $\nabla v(\mathbf{x})$ is almost surely non-zero.
Substituting these into equation (3.2.3), we obtain:
$\mathcal{L}_{\text{GSSM}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{\mathbf{A},\mathbf{b}}\mathbb{E}_{p_{d}}\left[\left((\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}s_{\theta}(\mathbf{X})\right)^{2}\right]$
$+\mathbb{E}_{\mathbf{A},\mathbf{b}}\mathbb{E}_{p_{d}}\left[(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})(\mathbf{A}\mathbf{X}+\mathbf{b})\right]$
$+\mathbb{E}_{\mathbf{A},\mathbf{b}}\mathbb{E}_{p_{d}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{A}(\mathbf{A}\mathbf{X}+\mathbf{b})\right]$
$+\mathbb{E}_{\mathbf{A},\mathbf{b}}\mathbb{E}_{p_{d}}\left[(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}s_{\theta}(\mathbf{X})\operatorname{tr}(\mathbf{A})\right]$ (54)
C.2 Variance Reduction for Quadratic GSSM
To reduce the variance of our estimator, we integrate out the randomness in $\mathbf{A}$ and $\mathbf{b}$. We compute the expectations of each term in the GSSM loss with respect to the randomness in $\mathbf{A}$ and $\mathbf{b}$, assuming $s_{\theta}(\mathbf{x})$ is independent of $\mathbf{A}$ and $\mathbf{b}$.
C.2.1 Expectation of the Quadratic Term
We first compute the expectation of the quadratic term in Equation (54):
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[\left((\mathbf{A}\mathbf{x}+\mathbf{b})^{\top}s_{\theta}(\mathbf{x})\right)^{2}\right]$ $=\mathbb{E}_{\mathbf{A}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)^{2}\right]$ (55)
$+2\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{b}\right)\right]+\mathbb{E}_{\mathbf{b}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{b}\right)^{2}\right]$ (56)
Since $\mathbf{A}$ and $\mathbf{b}$ have zero mean and are independent, the cross-term vanishes:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{b}\right)\right]=0$ (57)
Thus, we have:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[\left((\mathbf{A}\mathbf{x}+\mathbf{b})^{\top}s_{\theta}(\mathbf{x})\right)^{2}\right]=\mathbb{E}_{\mathbf{A}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)^{2}\right]+\mathbb{E}_{\mathbf{b}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{b}\right)^{2}\right]$ (58)
Computing $\mathbb{E}_{\mathbf{A}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)^{2}\right]$
We expand and compute:
$\mathbb{E}_{\mathbf{A}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)^{2}\right]$ $=\mathbb{E}_{\mathbf{A}}\left[\left(\sum_{i,j}s_{\theta i}(\mathbf{x})A_{ij}x_{j}\right)^{2}\right]$ (59)
$=\sum_{i,j,k,l}s_{\theta i}(\mathbf{x})s_{\theta k}(\mathbf{x})x_{j}x_{l}\mathbb{E}[A_{ij}A_{kl}]$ (60)
Using the properties of $\mathbf{A}$, the expectation $\mathbb{E}[A_{ij}A_{kl}]$ is non-zero only when $(i,j)=(k,l)$ or $(i,j)=(l,k)$. Therefore:
$\mathbb{E}[A_{ij}A_{kl}]=\begin{cases}\sigma_{1}^{2}&\text{if }i=j=k=l\\ \sigma_{2}^{2}&\text{if }i=k,j=l,i\neq j\\ \sigma_{2}^{2}&\text{if }i=l,j=k,i\neq j\\ 0&\text{otherwise}\end{cases}$ (61)
The expected value becomes:
$\mathbb{E}_{\mathbf{A}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\mathbf{x}\right)^{2}\right]$ $=\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{x})^{2}x_{i}^{2}+\sigma_{2}^{2}\left(\sum_{i\neq j}s_{\theta i}(\mathbf{x})^{2}x_{j}^{2}+\sum_{i<j}s_{\theta i}(\mathbf{x})s_{\theta j}(\mathbf{x})x_{i}x_{j}\right)$
$=\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{x})^{2}x_{i}^{2}+\sigma_{2}^{2}\left(\|\mathbf{x}\|^{2}\|s_{\theta}(\mathbf{x})\|^{2}+\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{x}\right)^{2}\right)$ (62)
Since the entries of $\mathbf{b}$ are independent with variance $\sigma_{3}^{2}$, we have:
$\mathbb{E}_{\mathbf{b}}\left[\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{b}\right)^{2}\right]=\sigma_{3}^{2}\|s_{\theta}(\mathbf{x})\|^{2}$ (63)
Combining both results, we obtain:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[\left((\mathbf{A}\mathbf{x}+\mathbf{b})^{\top}s_{\theta}(\mathbf{x})\right)^{2}\right]$ $=\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{x})^{2}x_{i}^{2}$
$+\sigma_{2}^{2}\left(\|\mathbf{x}\|^{2}\|s_{\theta}(\mathbf{x})\|^{2}+\left(s_{\theta}(\mathbf{x})^{\top}\mathbf{x}\right)^{2}\right)+\sigma_{3}^{2}\|s_{\theta}(\mathbf{x})\|^{2}$ (64)
C.2.2 Expectation of the Hessian Term
The Hessian term from Equation (54) is:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[s_{\theta}(\mathbf{x})^{\top}\mathbf{A}(\mathbf{A}\mathbf{x}+\mathbf{b})\right]=\mathbb{E}_{\mathbf{A}}\left[s_{\theta}(\mathbf{x})^{\top}\mathbf{A}^{2}\mathbf{x}\right]+\mathbb{E}_{\mathbf{A}}\left[s_{\theta}(\mathbf{x})^{\top}\mathbf{A}\right]\mathbb{E}_{\mathbf{b}}\left[\mathbf{b}\right]$ (65)
Since $\mathbb{E}_{\mathbf{A}}[\mathbf{A}]=\mathbf{0}$ and $\mathbb{E}_{\mathbf{b}}[\mathbf{b}]=\mathbf{0}$, the second term vanishes.
Computing $\mathbb{E}_{\mathbf{A}}\left[\mathbf{A}^{2}\right]$:
$\mathbb{E}_{\mathbf{A}}\left[\mathbf{A}^{2}\right]=\left(\sigma_{1}^{2}-\sigma_{2}^{2}\right)\mathbf{I}+n\sigma_{2}^{2}\mathbf{I}=\left(\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}\right)\mathbf{I}$ (66)
Therefore:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[s_{\theta}(\mathbf{x})^{\top}\mathbf{A}(\mathbf{A}\mathbf{x}+\mathbf{b})\right]=\left(\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}\right)s_{\theta}(\mathbf{x})^{\top}\mathbf{x}$ (67)
C.2.3 Expectation of the Trace Term
The trace term from Equation (54) is:
$\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[(\mathbf{A}\mathbf{x}+\mathbf{b})^{\top}s_{\theta}(\mathbf{x})\operatorname{tr}(\mathbf{A})\right]=\sigma_{1}^{2}s_{\theta}(\mathbf{x})^{\top}\mathbf{x}$ (68)
C.2.4 Final Variance-Reduced Loss
Combining Equations (64), (67), (68), and (54), we obtain the variance-reduced GSSM loss:
$\mathcal{L}_{\text{GSSM-VR}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{d}}\big{[}\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{X})^{2}X_{i}^{2}$ (69)
$+\sigma_{2}^{2}\left(\|\mathbf{X}\|^{2}\|s_{\theta}(\mathbf{X})\|^{2}+\left(s_{\theta}(\mathbf{X})^{\top}\mathbf{X}\right)^{2}\right)+\sigma_{3}^{2}\|s_{\theta}(\mathbf{X})\|^{2}\big{]}$
$+\mathbb{E}_{p_{d}}\left[\left(2\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}\right)s_{\theta}(\mathbf{X})^{\top}\mathbf{X}\right]$
$+\mathbb{E}_{p_{d}}\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})(\mathbf{A}\mathbf{X}+\mathbf{b})\right]$ (70)
C.3 Sampling from $\mathbf{A}\mathbf{x}$
In this section, we detail the construction of the random symmetric matrix $\mathbf{A}$ used in our quadratic transformation and describe how to efficiently sample $\mathbf{A}\mathbf{x}$ without explicitly forming $\mathbf{A}$.
Gaussian Orthogonal Ensemble (GOE) Matrices
We define $\mathbf{A}\in\mathbb{R}^{n\times n}$ as a random symmetric matrix drawn from the Gaussian Orthogonal Ensemble (GOE) with entries:
$A_{ii}$ $\sim\mathcal{N}(0,\sigma^{2}),\quad\text{for }i=1,\dots,n,$
$A_{ij}$ $=A_{ji}\sim\mathcal{N}(0,\sigma^{2}/2),\quad\text{for }i<j.$
This ensures that each off-diagonal element has variance $\sigma^{2}/2$, while diagonal elements have variance $\sigma^{2}$.
Distribution of $\mathbf{A}\mathbf{x}$
For a fixed vector $\mathbf{x}\in\mathbb{R}^{n}$, the product $\mathbf{A}\mathbf{x}$ is a zero-mean Gaussian vector:
$\mathbf{A}\mathbf{x}\sim\mathcal{N}\left(\mathbf{0},\,\boldsymbol{\Sigma}\right),$
where the covariance matrix $\boldsymbol{\Sigma}$ is given by:
$\boldsymbol{\Sigma}=\mathbb{E}\left[\mathbf{A}\mathbf{x}(\mathbf{A}\mathbf{x})^{\top}\right]=\frac{\sigma^{2}}{2}\left(\|\mathbf{x}\|^{2}\mathbf{I}_{n}+\mathbf{x}\mathbf{x}^{\top}\right).$
Efficient Sampling Procedure
To sample $\mathbf{A}\mathbf{x}$ without constructing $\mathbf{A}$, we use the following steps:
  1. 1.
    Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$.
  2. 2.
    Sample $z\sim\mathcal{N}(0,1)$, independently of $\boldsymbol{\epsilon}$.
  3. 3.
    Compute $\mathbf{A}\mathbf{x}$ as:
    $\mathbf{A}\mathbf{x}=\frac{\sigma}{\sqrt{2}}\left(\|\mathbf{x}\|\boldsymbol{\epsilon}+z\mathbf{x}\right).$
C.4 Additional Experimental Results
Beyond the score matching loss presented in the main text, we also evaluated the models using test log-likelihood. These estimates were computed using Annealed Importance Sampling (AIS) with a proposal distribution of $\mathcal{N}(0,2\mathbf{I})$ and $1,000,000$ samples, following the methodology of [23]. Figure 7 presents these results across all three UCI datasets.
Figure 7: Test log-likelihood for DKEF models trained with different loss functions. Higher values indicate better performance.
The test log-likelihood results largely align with our score matching loss findings. GSSM-VR demonstrates superior performance on RedWine and Parkinsons datasets, while showing slightly lower performance on WhiteWine.
Appendix D Deep Kernel Exponential Families with Generalized Sliced Score Matching
For our experiments, we follow the setup of [23], who demonstrated the effectiveness of sliced score matching with Deep Kernel Exponential Families (DKEFs). We maintain their model architecture and training procedure, changing only the loss function to our GSSM objective.
D.1 Deep Kernel Exponential Family Models
Following [27] and [23], we approximate the unnormalized log density as:
$\log\tilde{p}_{f}(\mathbf{x})=f(\mathbf{x})+\log q_{0}(\mathbf{x}),$ (71)
where $q_{0}(\mathbf{x})$ is a fixed function, and we represent $f$ as with Nyström approximation with $L$ inducing points $\{\mathbf{z}_{l}\}_{l=1}^{L}$:
$f(\mathbf{x})=\sum_{l=1}^{L}\alpha_{l}k(\mathbf{x},\mathbf{z}_{l}),$ (72)
where $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{L})^{\top}$ are the model parameters to be learned.
Following [23], we use their deep kernel architecture, which expresses the kernel as a mixture of 3 Guassian kernels:
$k_{w}(\mathbf{x},\mathbf{y})=\sum_{r=1}^{3}\rho_{r}\exp\left(-\frac{1}{2\sigma_{r}^{2}}\|\phi_{w_{r}}(\mathbf{x})-\phi_{w_{r}}(\mathbf{y})\|^{2}\right).$ (73)
In the above equation:
  • $\phi_{w_{r}}(\cdot)$ denotes a feature mapping associated with the $r$th component of the kernel, parameterized by $w_{r}$.
  • $\sigma_{r}$ represents the length scale for the $r$th component.
  • $\rho_{r}$ are nonnegative mixture coefficients that weight the contribution of each component to the overall kernel.
D.2 Quadratic Form of the GSSM Loss
As in standard score matching or sliced score matching, the GSSM loss also reduces to a quadratic form in terms of the model parameters $\boldsymbol{\alpha}$.
D.3 Quadratic Form of GSSM Loss
We now show that our GSSM loss, like SSM, becomes quadratic in $\boldsymbol{\alpha}$. For DKEF models, the score function is linear in $\boldsymbol{\alpha}$:
$s_{\boldsymbol{\theta}}(\mathbf{x})=\sum_{l=1}^{L}\alpha_{l}\nabla_{\mathbf{x}}k(\mathbf{x},\mathbf{z}_{l})+\nabla_{\mathbf{x}}\log q_{0}(\mathbf{x})$ (74)
From equation (3.2.3), our GSSM loss is:
$\mathcal{L}_{\text{GSSM}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\left(\nabla_{\mathbf{x}}v^{\top}(\mathbf{X})s_{\theta}(\mathbf{X})\right)^{2}\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\nabla_{\mathbf{x}}v^{\top}(\mathbf{X})\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\right]$
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\nabla_{\mathbf{x}}v^{\top}(\mathbf{X})s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right].$ (75)
Let us examine each term:
D.3.1 First Term
Define $g_{l}(\mathbf{x})=\nabla_{\mathbf{x}}k(\mathbf{x},\mathbf{z}_{l})$ and $h(\mathbf{x})=\nabla_{\mathbf{x}}\log q_{0}(\mathbf{x})$. Then:
$\left(\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})\right)^{2}$ $=\left(\nabla v(\mathbf{x})^{\top}\left(\sum_{l=1}^{L}\alpha_{l}g_{l}(\mathbf{x})+h(\mathbf{x})\right)\right)^{2}$
$=\left(\sum_{l=1}^{L}\alpha_{l}\nabla v(\mathbf{x})^{\top}g_{l}(\mathbf{x})+\nabla v(\mathbf{x})^{\top}h(\mathbf{x})\right)^{2}$ (76)
Define $\tilde{g}_{l}(\mathbf{x})=\nabla v(\mathbf{x})^{\top}g_{l}(\mathbf{x})$ and $\tilde{h}(\mathbf{x})=\nabla v(\mathbf{x})^{\top}h(\mathbf{x})$. Then:
$\frac{1}{2}\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\left(\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\right)^{2}\right]$ $=\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathbf{G}_{1}\boldsymbol{\alpha}+\boldsymbol{\alpha}^{\top}\mathbf{b}_{1}+C_{1}$ (77)
where
$(\mathbf{G}_{1})_{l,l^{\prime}}$ $=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\tilde{g}_{l}(\mathbf{X})\tilde{g}_{l^{\prime}}(\mathbf{X})\right]$ (78)
$(\mathbf{b}_{1})_{l}$ $=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\tilde{g}_{l}(\mathbf{X})\tilde{h}(\mathbf{X})\right]$ (79)
D.3.2 Second Term
We see:
$\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x})\nabla v(\mathbf{x})$ $=\nabla v(\mathbf{x})^{\top}\left(\sum_{l=1}^{L}\alpha_{l}\nabla_{\mathbf{x}}\nabla_{\mathbf{x}}^{\top}k(\mathbf{x},\mathbf{z}_{l})+\nabla_{\mathbf{x}}\nabla_{\mathbf{x}}^{\top}\log q_{0}(\mathbf{x})\right)\nabla v(\mathbf{x})$
$=\sum_{l=1}^{L}\alpha_{l}\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\nabla_{\mathbf{x}}^{\top}k(\mathbf{x},\mathbf{z}_{l})\nabla v(\mathbf{x})+\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\nabla_{\mathbf{x}}^{\top}\log q_{0}(\mathbf{x})\nabla v(\mathbf{x})$ (80)
Therefore:
$\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})\nabla v(\mathbf{X})\right]$ $=\boldsymbol{\alpha}^{\top}\mathbf{b}_{2}+C_{2}$ (81)
where
$(\mathbf{b}_{2})_{l}$ $=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\nabla v(\mathbf{X})^{\top}\nabla_{\mathbf{x}}\nabla_{\mathbf{x}}^{\top}k(\mathbf{X},\mathbf{z}_{l})\nabla v(\mathbf{X})\right]$ (82)
D.3.3 Third Term
For the Hessian term:
$s_{\theta}(\mathbf{x})^{\top}\mathbf{H}_{v}(\mathbf{x})\nabla v(\mathbf{x})$ $=\left(\sum_{l=1}^{L}\alpha_{l}g_{l}(\mathbf{x})+h(\mathbf{x})\right)^{\top}\mathbf{H}_{v}(\mathbf{x})\nabla v(\mathbf{x})$
$=\sum_{l=1}^{L}\alpha_{l}g_{l}(\mathbf{x})^{\top}\mathbf{H}_{v}(\mathbf{x})\nabla v(\mathbf{x})+h(\mathbf{x})^{\top}\mathbf{H}_{v}(\mathbf{x})\nabla v(\mathbf{x})$ (83)
Therefore:
$\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla v(\mathbf{X})\right]$ $=\boldsymbol{\alpha}^{\top}\mathbf{b}_{3}+C_{3}$ (84)
where
$(\mathbf{b}_{3})_{l}$ $=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[g_{l}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla v(\mathbf{X})\right]$ (85)
D.3.4 Fourth Term
For the Laplacian term:
$\nabla v(\mathbf{x})^{\top}s_{\theta}(\mathbf{x})\Delta v(\mathbf{x})$ $=\Delta v(\mathbf{x})\left(\sum_{l=1}^{L}\alpha_{l}\nabla v(\mathbf{x})^{\top}g_{l}(\mathbf{x})+\nabla v(\mathbf{x})^{\top}h(\mathbf{x})\right).$ (86)
Therefore:
$\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\nabla v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right]$ $=\boldsymbol{\alpha}^{\top}\mathbf{b}_{4}+C_{4}$ (87)
where
$(\mathbf{b}_{4})_{l}$ $=\mathbb{E}_{p_{v}}\mathbb{E}_{p_{d}}\left[\Delta v(\mathbf{X})\nabla v(\mathbf{X})^{\top}g_{l}(\mathbf{X})\right]$ (88)
D.3.5 Complete Loss
Combining all terms, the total loss is quadratic in $\boldsymbol{\alpha}$:
$\mathcal{L}_{\text{GSSM}}(\boldsymbol{\alpha})=\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathbf{G}_{1}\boldsymbol{\alpha}+\boldsymbol{\alpha}^{\top}(\mathbf{b}_{1}+\mathbf{b}_{2}+\mathbf{b}_{3}+\mathbf{b}_{4})+C$ (89)
As in [23], we add $\ell_{2}$ regularization:
^$\hat{\mathcal{L}}_{\text{GSSM}}(\boldsymbol{\alpha})=\mathcal{L}_{\text{GSSM}}(\boldsymbol{\alpha})+\frac{1}{2}\lambda_{\boldsymbol{\alpha}}\|\boldsymbol{\alpha}\|^{2}$ (90)
yielding the solution:
$\boldsymbol{\alpha}^{*}=-(\mathbf{G}_{1}+\lambda_{\boldsymbol{\alpha}}\mathbf{I})^{-1}(\mathbf{b}_{1}+\mathbf{b}_{2}+\mathbf{b}_{3}+\mathbf{b}_{4})$ (91)
The derivation for the variance-reduced version follows similarly.
D.4 Implementation Details
All hyperparameters (such as network architectures, optimization settings, and the choice of $q_{0}(\mathbf{x})$) follow [23] to ensure a fair comparison and reproducibility. The only difference is that we replace the original loss with our GSSM-based loss. This maintains consistency with previous experiments while highlighting the direct impact of our new score matching objective.
Appendix E Detailed Derivations for Chess Diffusion
E.1 SDE Parameters and Transformation
Let $\mathbf{f}(\mathbf{x},t)$ and $g(t)$ be the drift and diffusion coefficients. We use a Variance Preserving (VP) SDE with:
$\mathbf{f}(\mathbf{x},t)$ $=-\frac{1}{2}\beta(t)\mathbf{x},$ (92)
$g(t)$ $=\sqrt{\beta(t)},$ (93)
where $\beta(t)$ is the noise schedule. The additive logistic transformation is:
$\phi_{i}(\mathbf{x})$ $=\frac{e^{x_{i}}}{1+\sum_{j=1}^{12}e^{x_{j}}},\quad i=1,\dots,12,$ (94)
which maps vectors in $\mathbb{R}^{12}$ to the probability simplex.
E.2 Transformed Drift and Diffusion Coefficients
The transformed drift $\bar{\mathbf{f}}(\mathbf{y},t)$ and diffusion $\bar{\mathbf{G}}(\mathbf{y},t)$ are:
$\bar{f}_{i}(\mathbf{y},t)$ $=y_{i}\left[k_{i}-\sum_{j=1}^{12}y_{j}k_{j}\right],$ (95)
$\bar{G}_{ij}(\mathbf{y},t)$ =$=\begin{cases}g(t)y_{i}(1-y_{i}),&\text{if }i=j,\\ -g(t)y_{i}y_{j},&\text{if }i\neq j,\end{cases}$ (96)
where $k_{i}=f_{i}(\mathbf{x},t)+\frac{1}{2}g^{2}(t)(1-2y_{i})$.
E.3 Score Function Transformation
The score function $\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y})$ in terms of $\mathbf{x}$ is:
$(\nabla_{\mathbf{y}}\log q_{t}(\mathbf{y}))_{i}$ $=\left(\mathbf{J}_{\phi^{-1}}(\mathbf{y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\ \mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x},t),t)^{\top}\right)$ (97)
$=\frac{1}{y_{13}}\sum_{j=1}^{12}\frac{\partial}{\partial x_{j}}\log p_{t}(\mathbf{x})+\frac{1}{y_{i}}\frac{\partial}{\partial x_{i}}\log p_{t}(\mathbf{x})+\frac{y_{i}-y_{13}}{y_{i}y_{13}},$ (98)
where $y_{13}=1-\sum_{j=1}^{12}y_{j}$.
E.4 Divergence Term
The divergence term $\nabla_{\mathbf{y}}\cdot\left(\bar{\mathbf{G}}(\mathbf{y},t)\bar{\mathbf{G}}(\mathbf{y},t)^{\top}\right)$ for the $i$-th component is:
$(\nabla_{\mathbf{y}}\cdot\left(\bar{\mathbf{G}}(\mathbf{y},t)\bar{\mathbf{G}}(\mathbf{y},t)^{\top}\right))_{i}$ $=g(t)^{2}y_{i}\left(2-2S_{y}+15S_{y^{2}}-15y_{i}\right),$ (99)
where $S_{y}=\sum_{j=1}^{12}y_{j}$ and $S_{y^{2}}=\sum_{j=1}^{12}y_{j}^{2}$.
E.5 Training Details
We trained a score-based model $s_{\theta}(\mathbf{x},t)$ using a standard U-Net architecture [22]. The model was trained on a dataset of chess positions obtained from Kaggle [5]. Training was conducted using an NVIDIA GeForce RTX 4090 GPU.